### v3

# DualPM: Dual Posed-Canonical Point Maps for 3D Shape and Pose Reconstruction

Abstract
The choice of data representation is a key factor in the success of deep learning in geometric tasks. For instance, DUSt3R has recently introduced the concept of viewpoint-invariant point maps, generalizing depth prediction, and showing that one can reduce all the key problems in the 3D reconstruction of static scenes to predicting such point maps. In this paper, we develop an analogous concept for a very different problem, namely, the reconstruction of the 3D shape and pose of deformable objects. To this end, we introduce the Dual Point Maps (DualPM), where a pair of point maps is extracted from the same image, one associating pixels to their 3D locations on the object, and the other to a canonical version of the object at rest pose. We also extend point maps to amodal reconstruction, seeing through self-occlusions to obtain the complete shape of the object. We show that 3D reconstruction and 3D pose estimation reduce to the prediction of the DualPMs. We demonstrate empirically that this representation is a good target for a deep network to predict; specifically, we consider modeling horses, showing that DualPMs can be trained purely on 3D synthetic data, consisting of a single model of a horse, while generalizing very well to real images. With this, we improve by a large margin previous methods for the 3D analysis and reconstruction of this type of objects.
Figure 1: Left: We map an image of an object to its Dual Point Maps (DualPMs), a pair of point maps $P$, defined in a camera space, and $Q$, defined in a canonical space where the object has a neutral pose. The pose is thus given by the flow $P-Q$. Right: The DualPMs are easy to predict with a neural network, enabling effective 3D object reconstruction and facilitating geometric tasks like detecting 3D keypoints and fitting a 3D skeleton. For visualization, we color each point with its coordinate in the canonical point maps.
**footnotetext: Equal contribution.
1 Introduction
An important question in 3D computer vision is finding the optimal way of interfacing visual geometry and neural networks. One approach is to use neural networks for pre-processing, for instance, to detect and match image keypoints. Once these 2D primitives are extracted, visual geometry can be used to infer the 3D structure of the scene by solving a system of equations or via optimization [62]. The alternative approach is to task the neural network with outputting directly the geometric quantities of interest, such as depth maps [77] or the camera pose [61].
Until recently, the general consensus was that inferring accurate 3D information should be, whenever possible, left to visual geometry. Hence, neural networks would fill gaps such as feature matching and monocular prediction where visual geometry is not applicable. However, this view is increasingly challenged as authors find better ways of reducing geometric tasks to the calculation of quantities that can be predicted very effectively by neural networks. Most recently, DUSt3R [65, 8] has demonstrated the power of predicting point maps, namely images that associate to each pixel its corresponding 3D point in the scene. They show that many tasks in the reconstruction of static scenes, such as matching, camera estimation, and triangulation, can be solved trivially from the point maps predicted by a neural network.
In this work, we ask whether a similar intuition is applicable to a very different class of problems in visual geometry, namely the monocular reconstruction of the 3D shape and pose of deformable objects. We do so by developing a new network-friendly representation, which we call Dual Point Maps (DualPMs).
Our setting is a substantial departure from DUSt3R, and so are their point maps and our DualPMs. Nevertheless, the starting point is the same: a point map in both DUSt3R and DualPM associates to each image pixel its corresponding 3D point on the object. Predicting this point map gives us a 3D reconstruction of the object, which is useful, but gives us no information about the pose of the object.
To clarify this further, consider the example of a horse. Its pose is defined as the deformation that takes the horse from its neutral pose to the pose observed in the image. Hence, reconstructing its pose means finding the transformation, or deformation field, between two versions of the 3D object: the one seen in the image and the canonical version of the same, in neutral pose. Knowledge of a single point map gives no information about this transformation. So the question is: How can we extend point maps so that recovering this deformation field becomes trivial?
To solve this problem, our DualPMs predicts not one, but two point maps from the same image (Fig. 1). They map each pixel to two versions of the same 3D point. The first version is the usual reconstruction of the object in 3D. The second version, however, is the location of the point in the canonical version of the object. With this representation, it is easy to recover the pose of the object: in fact, the deformation field that we wish to recover is simply the difference between the two point maps. Knowledge of this deformation field is in turn useful for various analysis tasks, such as recovering 3D keypoints or estimating the object’s articulation in terms of a skeleton (see  Fig. 1).
Figure 2: An amodal point map associates each pixel $\boldsymbol{u}$ to an even number of objects points $\boldsymbol{p}_{i}$, corresponding to the locations where the pixel’s ray intersects the object’s surface. Predicting an amodal point map reconstructs the entire object despite self-occlusions.
The canonical version of the point map is also viewpoint invariant and bears some resemblance to the Normalized Object Coordinates of [59] and the Canonical Surface Maps of [56, 24]. As such, its prediction is similar to a keypoint labeling problem, which can be solved effectively by a neural network. Furthermore, we show that the canonical point map is also a good feature map in its own right; in fact, our model predicts the posed point map from the canonical one, using it as a feature extractor instead of, for instance, DINOv2 [44].
A shortcoming of point maps, in general, is that they only reconstruct the visible part of the object. To address this issue, we further introduce the idea of estimating amodal point maps; namely, each pixel is mapped not only to the visible point of the object but to all points on the object that project onto that pixel, regardless of self-occlusions (Fig. 2). To effectively represent such point maps for neural network prediction, we propose a layered representation. In this representation, the first layer encodes the visible points, and each subsequent layer captures the next set of points occluded by the previous layer. This concept is similar to depth peeling, a standard technique used in rasterization in computer graphics. With this extension, we can provide a full reconstruction of the object’s shape and pose.
To test our ideas, we consider the problem of reconstructing the 3D shape and pose of horses from monocular images. We do so as horses represent a widely studied category of deformable objects [25, 69, 71, 17, 32], with synthetic 3D data readily available for this class [17], enabling controlled evaluations. Furthermore, we show that synthetic data derived from a single horse model is sufficient to train a DualPM model hat generalizes robustly to real-world data thanks to our effective representation that makes learning relatively easy. With this, we show that we can improve previous methods by a large margin for the 3D analysis and reconstruction of this type of objects, even on of out-of-domain real-world datasets, both in terms of correspondence and 3D reconstruction. Moreover, our method exhibits strong zero-shot generalization to real-world images of other quadruped animals, such as cow and sheep, despite being trained exclusively on a single horse model. We also demonstrate the ability of the point maps to reconstruct the pose of the object in terms of their deformation fields, and suggest applications such as fitting articulated 3D skeletons to the object that can be used for motion transfer and animations.
To summarize, our contributions are as follows:
  1. 1.
    We introduce the novel concept of DualPM, which reduces monocular 3D shape and pose reconstruction to the prediction of a pair of point maps from the image: one defined in canonical space and the other in the posed camera space.
  2. 2.
    We extend point maps to enable amodal reconstruction along with their efficient layered representation, producing the complete 3D shape of the articulated object.
  3. 3.
    We demonstrate that DualPMs can be effectively predicted by a neural network and that this representation serves as a suitable target for learning, to the extent that synthetic data is sufficient for training the model.
  4. 4.
    We show key tasks in the analysis of deformable objects, including 3D reconstruction, keypoint transfer/matching, estimation of deformation, and fitting of articulated skeletons, reduce to the prediction of DualPMs.
  5. 5.
    We demonstrate that DualPMs significantly outperform prior approaches for 3D analysis and reconstruction of horses and generalize effectively to other quadrupeds, even when trained exclusively on synthetic data from a single articulated horse model.
2 Related work
Deformable reconstruction.
Our work is related to the reconstruction of deforming scenes. Many such works consider the optimization approach, where a 4D model of the scene is fitted to a video sequence. In works like Video Pop-up[50], Monocular 3D [26], MonoRec [66] the 3D reconstruction of different video frames are regularized, for instance to enforce approximate local rigidity.
More often, the scene 4D scene is represented as the deformation of a canonical configuration. DynamicFusion [42] is an early example of this approach. More recent works like Neural Volumes [39] reconstruct both shape and appearance. D-NeRF [46], Neural Radiance Flow [7], Neural Scene Flow Fields [30], Dynamic Video Synthesis [15], DynIBaR [31], MorpheuS [60], extend NeRF [41] with dense deformation fields to capture motion. Similarly, 4D Gaussian Splatting [67], Gaussian Flow [35], DynMF [23], GauFRe [33], MoSca [27], Street Gaussians [73], Dynamic Gaussian Marbles [55], Shape of Motion [64], Ego Gaussians [83], 4DGS [67] also estimate deformations, but in the context of 3D Gaussian Splatting [20] rather than NeRF.
Rather than modelling dense deformation fields, approaches like Dynamic NeRF [14], Seeing 3D Objects [51], K-planes [13] and HexPlane [4] just add time or pose parameterization to the NeRF model. Approaches like 4D Visualization [2] and NVSD [82] predict novel views of a dynamic scene without explicit 3D reconstruction.
Fourier PlenOctrees [63], NeRFPlayer [53] and RT-4DGS [78] focus on the problem of storing and rendering 4D models efficiently. 4DGen [81], Align Your Gaussians [36], DreamScene4D [6], L4GM [47] considered instead the problem of generating 4D models from text (text-to-4D). PhysGaussian [72], MD-Splatting [9], Gaussian Splashing [12] and VR-GS [18] consider reconstruction and generation of 4D scene which accounts for physical principles
Works like Neural Human Video Rendering [38], AutoAvatar [1], Neural Body [45], Dynamic Facial RF [52], Relighting4D [5], Animate124 [86], Dynamic Gaussian Mesh [37] and IM4D [34] specialise in the reconstruction of articulated characters like humans and animals. Often these define the object in a canonical neutral pose, and encode pose as a deformation applied to this pose.
Learning deformable 3D object categories.
Works like 3D Menagerie [87], Unsupervised 3D [68], LASR [74], BANMo [75], Dessie [28], DOVE [69], LASSIE [79], MagicPony [70], Farm3D [17], 3D Fauna [32], and others [16, 88, 76, 80, 49] have considered the problem of learning models of 3D object categories. In these models, pose is expressed as the object deformation with respect to a version of the object in neutral pose, which usually takes the form of a category-level template. Such a template can be learned either in a weakly-supervised manner [69, 79, 70, 17, 32, 76, 80] or from 3D data, most notably SMPL [40], SMAL [3], and more recently, VAREN [89]. Concurrent Dessie also introduces a synthetic data generation pipeline similar to ours.
Point maps and canonical maps.
DUSt3R [65] was the first paper to intuit the power of predicting point maps instead of depth for 3D reconstruction. The follow up MASt3R [8] builds on this to obtain a state-of-the-art Structure-from-Motion (SfM) system. MonST3R [85] extend DUSt3R to monocular video, reconstructing dynamic point maps in a fixed reference frame.
The concept of learning a map that sends pixels to canonical object points can be found in the dense equivariant mapping of [56, 57, 58]. Normalized object coordinates NOCs [59] modify this idea for rigid objects by sending pixels to 3D points defined in the object’s surface in canonical pose. Our approach also does this, but with two key differences. First, the input object is deformable rather than rigid. Second, we predict dual point maps, which encode the object deformation beyond a simple rigid transformation.
Figure 3: Method overview. We preprocess the input image $I$ by obtaining a segmentation mask $M$ and image features $F$ using pretrained networks. Subsequently, we predict the layered canonical point map $\{Q_{i}^{*}\}_{i=1}^{2K}$. Conditioned on $\{Q_{i}^{*}\}_{i=1}^{2K}$, we predict the layered posed point map $\{P_{i}^{*}\}_{i=1}^{2K}$, along with the associated opacity maps $\{\sigma_{i}^{*}\}_{i=3}^{2K}$. Both layered point maps are then decoded to produce the canonical point map $Q$ and the posed point map $P$. The training process is supervised using the ground truth point maps and opacity maps.
3 Method
Let $I\in\mathbb{R}^{3\times H\times W}$ be an image containing a deformable object we wish to reconstruct in 3D, along with its pose. We assume that a mask $M\in\{0,1\}^{H\times W}$ is available, indicating the object’s location in the image. Next, we introduce the concept of dual point maps (Sec. 3.1) and then of its amodal extension (Sec. 3.2).
3.1 Dual point maps
A point map $P\in\mathbb{R}^{3\times H\times W}$ is a tensor containing a 3D point for each pixel of the input image $I$. We are, in particular, interested in the pixels $\boldsymbol{u}\in M$ contained within the object’s mask $M$. Then $\boldsymbol{u}$ is the image of some point $\boldsymbol{p}$ that belongs to the surface of the object, known as the imaged point. The image sample $I(\boldsymbol{u})$ is the color of the imaged point (an RGB triplet), and $\boldsymbol{p}=P(\boldsymbol{u})\in\mathbb{R}^{3}$ is the 3D location of the point, which we express in the reference frame of the camera.
By estimating the point map $P$ from the image $I$, we reconstruct the visible portion of the 3D object. This is similar to predicting a depth map but provides strictly more information. In fact, the 3D point does not depend on the camera’s intrinsic parameters, such as the focal length, as it is directly expressed in the camera’s reference frame. As shown in [65], this fact can be used to infer the camera intrinsics. However, this point map does not capture the object’s deformation and thus cannot be used by itself to recover its pose.
We address this limitation by introducing the novel concept of Dual Point Maps (DualPMs). In addition to $P$, we predict a second point map $Q\in\mathbb{R}^{3\times H\times W}$ defined in canonical space. By canonical, we mean two things: first, the points in $Q$ are expressed in the object reference frame rather than in the camera one; second, the object is in a canonical pose, usually a neutral pose. Hence, differently from $P$, the point map $Q$ is invariant to the object’s pose and deformation.
The pair of point maps $P$ and $Q$ implicitly encode the object deformation: they tell us that pixel $\boldsymbol{u}$ corresponds to a point which is found at location $\boldsymbol{q}=Q(\boldsymbol{u})$ in rest pose and at location $\boldsymbol{p}=P(\boldsymbol{u})$ once posed. This provides rich information about the object pose. For example, given two images $I_{1}$ and $I_{2}$ of the object in two different poses, we can match a pixel $\boldsymbol{u}_{1}$ in the first image to its corresponding pixel $\boldsymbol{u}_{2}$ in the second by matching the corresponding canonical points $Q_{1}(\boldsymbol{u}_{1})$ and $Q(\boldsymbol{u}_{2})$, i.e.,
$\boldsymbol{u}_{2}(\boldsymbol{u}_{1})=\operatornamewithlimits{argmin}_{\boldsymbol{u}_{2}\in M_{2}}\|Q_{1}(\boldsymbol{u}_{1})-Q_{2}(\boldsymbol{u}_{2})\|^{2}.$
This works because the canonical points are view and pose invariant. Then, we can infer the so-called ‘scene flow’, or motion between the 3D points, as $P_{2}(\boldsymbol{u}_{2}(\boldsymbol{u}_{1}))-P_{1}(\boldsymbol{u}_{1}).$
Predicting the dual point maps.
To predict the dual point maps from the image $I$, we introduce a predictor neural network $(P,Q)=\Phi(I).$ Its architecture is illustrated in Fig. 3. A key property of the canonical point map $Q$ is its invariance to the object’s pose and deformation. This invariance simplifies its prediction by a neural network, as it reduces it to a pixel labeling problem. To further ease the prediction of $Q$ and improve generalization, we leverage recent advances in self-supervised image feature learning [84, 44, 10], and condition its prediction on strong features $F$ extracted from the image $I$ using a pre-trained network $\Psi$, $F=\Psi(I)$ from [84]. These features serve as a good proxy for the canonical point map, as they have been shown to be nearly invariant to the object’s pose and deformation, thereby simplifying the network’s task. We then condition the prediction of $P$ on $Q$ instead of $F$ as it improves generalization of the model on out-of-distribution images, as demonstrated in Sec. 4.5. This shows that learning canonical point maps as an intermediate representation benefits the 3D reconstruction task. To summarize, we first predict $Q$ using the features $F$, and then condition the prediction of $P$ on $Q$:
$Q=\Phi_{Q}(\Psi(I)),\quad P=\Phi_{P}(Q).$
We extend both networks $\Phi_{Q}$ and $\Phi_{P}$ to also predict a per-pixel confidence score. So, for example, the map $\Phi_{P}$ outputs the map $c_{P}(\boldsymbol{u})>0$, which is used to train the model by means of the self-calibrated L2 loss of [43, 19, 65]:
$\mathcal{L}_{P}=\frac{1}{|M|}\sum_{\boldsymbol{u}\in M}c_{P}(\boldsymbol{u})\|\hat{P}(\boldsymbol{u})-P(\boldsymbol{u})\|^{2}-\alpha\log c_{P}(\boldsymbol{u}),$
where $\alpha>0$ is a constant and $\hat{P}$ is the ground-truth point map. The analogous loss $\mathcal{L}_{Q}$ is used for $\Phi_{Q}$.
3.2 Amodal point maps
A limitation of the dual point maps is that they only capture the visible portion of the object. We now show how to predict a complete 3D point cloud for the object, including the points that are not visible due to self-occlusion.
To this end, we introduce the concept of an amodal point map. Namely, we associate each pixel $\boldsymbol{u}\in M$ within the object mask to the sequence of 3D points $(\boldsymbol{p}_{1},\boldsymbol{p}_{2},\dots)$ that intersect the camera ray through $\boldsymbol{u}$ in order of increasing distance from the camera center. Hence, $\boldsymbol{p}=\boldsymbol{p}_{1}$ is the imaged point considered in Sec. 3.1, $\boldsymbol{p}_{2}$ is the point that would be imaged should we pluck a hole in the object surface at $\boldsymbol{p}_{1}$, and so on. Then, the goal is to predict the amodal point map $\mathcal{P}$, defined as the map that associates to each pixel $\boldsymbol{u}$ the sequence of 3D points $\mathcal{P}(\boldsymbol{u})=(\boldsymbol{p}_{1},\boldsymbol{p}_{2},\dots)$.
Note that the length of the sequence $\mathcal{P}(\boldsymbol{u})$ is not fixed but depends on the pixel $\boldsymbol{u}$. The sequence length is always even because the ray must exit the object once it enters it, as we assume the camera is positioned outside the object. To account for this, we extend the network $\Phi_{P}$ by first predicting a pair of points $(\boldsymbol{p}_{1},\boldsymbol{p}_{2})_{1}$ for each pixel $\boldsymbol{u}$. Subsequently, additional sets of outputs $(\boldsymbol{p}_{2k-1},\boldsymbol{p}_{2k})_{k}$ are added for $k\in[2,K]$, where $K$ denotes the total number of pairs, capturing further potential ray intersections with the object. Unlike the first set of intersections, these additional intersections may not always occur; therefore, we also predict opacities $(\sigma_{2k-1},\sigma_{k})_{k}$ for these cases. As a result, each pixel $\boldsymbol{u}$ is mapped to $3\times 2+(3+1)\times 2\times(K-1)$ scalar outputs. The network $\Phi_{Q}$ is extended in a similar manner. The losses $\mathcal{L}_{P}$ and $\mathcal{L}_{Q}$ are also extended to account for the amodal point maps. We also add additional loss $\mathcal{L}_{\sigma}$ equivalent to those of $\mathcal{L}_{P}$ and $\mathcal{L}_{Q}$ to supervised the opacities prediction.
Layered point maps.
To efficiently represent the amodal point map for the neural network prediction, we introduce an image-based layered representation for DualPMs denoted as $(P^{*},Q^{*})$. In this representation, the first layer encodes the visible points, and each consecutive layer captures the next set of points occluded by the previous layer. We also represent the opacity maps $\sigma^{*}$ in the same layered fashion.
Since we supervise our method with synthetic data, we can easily generate training targets for this representation. Each vertex of the object mesh is assigned a 6D attribute, consisting of the calculated posed vertex 3D position and the corresponding vertex 3D position in canonical space. The mesh is then rendered from the camera’s viewpoint. During the rendering process, the rasterizer generates a set of pixel-aligned layers, $L_{i}\in\mathbb{R}^{6\times H\times W}$, which contain the rendered 6D attributes, together with the associated opacity map $\hat{\sigma}_{i}^{*}\in[0,1]^{H\times W}$ for every ray-object intersection. We iterate through the rasterized layers from the front to the back until the number of layers specified for our model is reached. This approach yields ground-truth DualPMs with a number of layers corresponding to the ray-object intersections. We split the attributes of each layer corresponding to the canonical and posed positions, yielding ground-truth layered canonical point maps
$\hat{Q}^{*}=(\hat{Q}_{1}^{*},\hat{Q}_{2}^{*},\dots,\hat{Q}_{N}^{*}),\quad\hat{Q}_{i}^{*}\in\mathbb{R}^{3\times H\times W},$
and layered posed point maps
$\hat{P}^{*}=(\hat{P}_{1}^{*},\hat{P}_{2}^{*},\dots,\hat{P}_{N}^{*}),\quad\hat{P}_{i}^{*}\in\mathbb{R}^{3\times H\times W},$
representations with their associated opacity maps $\sigma_{i}^{*}$. The number of layers $N$ corresponds to $2K$. The canonical $Q$ and posed $P$ point maps can be then simply extracted from this representation by masking out transparent points.
3.3 Training
Figure 4: Synthetic training data. We generate synthetic training data by rendering a single rigged 3D horse model with two texture variations (top row) in various poses with different environmental maps and under random viewpoints (bottom row).
We train our model using synthetic data generated from a single articulated 3D horse model sourced from the Animodel dataset [17]. Following the data generation pipeline described in Animodel, we randomly sample object poses, camera poses, and illumination conditions. These are modeled by randomly sampling from collection environment maps. For each sample, we render the object into an image $I$ and generate the corresponding image-based layered paired point maps $P^{*}$ and $Q^{*}$, which serve as targets for the model, using the procedure detailed in Sec. 3.2. We illustrate our synthetic training data in Fig. 4.
4 Experiments
3D-Fauna [32]
Ours
Figure 5: Comparison with the state-of-the-art. We compare our method with the state-of-the-art method 3D-Fauna [32]. For visualization, we paint each point of the recovered point map $P$ with its canonical coordinate. Our reconstructed shape aligns more closely with the object’s structure compared to 3D-Fauna and adapts well to large variations within the object category, despite being trained on a single adult male horse model. Our method reliably recovers shapes outside its training distribution, such as foals and dwarf horses (3rd and 4th rows from the bottom). It also performs well on very challenging poses, such as those shown in the last two rows, where 3D-Fauna completely fails.
Method PCK (%) Chamfer Distance (cm)
Horse Cow Sheep Real-Sized Normalized
Horse Cow Sheep Horse Cow Sheep
A-CSM [25] 32.9 26.3 28.6 7.60 ± $\pm\ 3.07$ 6.71 ± $\pm\ 1.81$ 2.84 ± $\pm\ 0.77$ 2.73 ± $\pm\ 1.13$ 2.35 ± $\pm\ 0.68$ 2.48 ± $\pm\ 0.70$
MagicPony [71] 42.9 42.5 41.2 7.19 ± $\pm\ 2.35$ 7.22 ± $\pm\ 1.53$ 3.43 ± $\pm\ 0.73$ 2.58 ± $\pm\ 0.80$ 2.53 ± $\pm\ 0.59$ 3.00 ± $\pm\ 0.68$
Farm3D [17] 49.1 40.2 36.1 7.65 ± $\pm\ 2.21$ 6.91 ± $\pm\ 1.49$ 3.79 ± $\pm\ 0.55$ 2.76 ± $\pm\ 0.83$ 2.41 ± $\pm\ 0.54$ 3.31 ± $\pm\ 0.49$
3D-Fauna [32] 53.9 8.69 ± $\pm\ 2.38$ 9.19 ± $\pm\ 2.40$ 3.51 ± $\pm\ 0.88$ 3.13 ± $\pm\ 0.85$ 3.20 ± $\pm\ 0.80$ 3.06 ± $\pm\ 0.76$
Ours 73.2 63.0 64.2 3.13 ± $\pm\ 2.13$ 4.74 ± $\pm\ 1.40$ 2.32 ± $\pm\ 0.78$ 1.11 ± $\pm\ 0.73$ 1.67 ± $\pm\ 0.55$ 2.03 ± $\pm\ 0.71$
Table 1: Quantitative evaluation. We evaluate on PASCAL VOC, reporting PCK@0.1 (higher is better $\uparrow$), and on Animodel [17], reporting the bi-directional Chamfer Distance in centimeters (lower is better $\downarrow$). Our model, trained solely on data from a single horse model, outperforms state-of-the-art approaches trained on each of the respective categories.
We evaluate our DualPM in terms of both qualitative (Sec. 4.1) and quantitative performance (Sec. 4.2) on the horse category, demonstrate zero-shot generalization to unseen categories of quadrupeds (Fig. 6), show that our DualPM can be used to fit 3D skeletons for motion transfer and animations (Sec. 4.4), and conduct an ablation study to assess the impact of our design choices on model performance (Sec. 4.5). We include technical and implementation details in App. A.
4.1 Qualitative evaluation
We qualitatively evaluate our method on challenging in-the-wild images of horses from the PASCAL VOC dataset [11]. The results are presented in Figs. 1 and 5. Our method successfully recovers the 3D shape and pose of horses from a single image, despite wide variations in pose, shape, and appearance within the horse category, including foals and dwarf horses, even though it is trained on a single adult horse model. We also compare our method with the state-of-the-art approach for single-view reconstruction of deformable objects, 3D-Fauna [32], and present side-by-side comparisons in Fig. 5. Our method performs better in recovering the shape and pose than 3D-Fauna, which struggles to follow the object’s structure accurately and fails to reconstruct intricate details specific to each object.
4.2 Quantitative evaluation
Following prior work on deformable object reconstruction [25, 29, 71, 32], we evaluate our method on the keypoint transfer task using PASCAL VOC dataset [11]. We adhere to the protocol of [25] and report the percentage of correctly transferred keypoints with a 10% threshold (PCK@0.1) between pairs of images. The results are presented in Tab. 1, where we compare our method with state-of-the-art approaches.
We also measure the bi-directional Chamfer distance between the predicted point maps and the ground truth 3D shape on the Animodel dataset introduced in Farm3D [17] . This dataset comprises renderings of 3D models of animals, including horses, cows, and sheep, in various poses, representing the shapes of both male and female specimens. As our model is trained using renderings of the male horse from this dataset, we evaluate its performance, along with other approaches, on a subset containing only the female horse. To ensure a fair comparison, we also exclude the poses used in this benchmark from our training set. The performance of our method and others is reported in Tab. 1.
Our method consistently outperforms all other approaches on both benchmarks, demonstrating its ability to generalize to unseen poses and shapes. Moreover, it exhibits exceptional generalization on real-world images despite being trained solely on synthetic data.
4.3 Generalization to unseen categories
Figure 6: Results on unseen categories. Our method demonstrates robust generalization to the unseen categories such as cow and sheep, despite being trained solely with a single horse model.
Given the generalization capabilities of our method demonstrated on the horse category, we extend our analysis to unseen quadruped categories, specifically cows and sheep. We evaluate our approach using the same datasets as for the horses, PASCAL VOC [11] and Animodel [17], following the same evaluation protocol, with results reported in Tab. 1. Furthermore, we provide additional qualitative results on the same dataset in Fig. 6. Our method exhibits strong zero-shot generalization to these categories, outperforming state-of-the-art approaches on both datasets, despite being trained exclusively on a single horse model.
4.4 Skeleton fitting and animation
Figure 7: Animation. DualPMs allow for fitting a 3D skeleton, which can subsequently be used to animate the 3D reconstruction by utilizing existing animations through motion retargeting.
With a predicted DualPM $(P,Q)$, we can easily fit an instance-specific 3D skeleton. Our method is trained on synthetic data from a model rigged with a skeleton composed of bones $B$ that deform the mesh. These bones are defined in the canonical space, and each bone is associated with a set of vertices on the mesh through skinning weights.
For each point $(\boldsymbol{p}_{i},\boldsymbol{q}_{i})$ of the DualPM $(P,Q)$, we identify the closest vertex of the mesh in the canonical space to the canonical point $\boldsymbol{q}_{i}$ then we associate the point $(\boldsymbol{p}_{i},\boldsymbol{q}_{i})$ with the skinning weights of that vertex. For each bone $b\in B$, we identify subsets of $\{(\boldsymbol{p}_{i},\boldsymbol{q}_{i})_{b}\}$, where the associated skinning weights between the bone $b$ and the points $(\boldsymbol{p}_{i},\boldsymbol{q}_{i})$ are higher than 0.5. We then compute the transformation that aligns the subset of canonical points $\{(\boldsymbol{q}_{i})_{b}\}$ to the subset of posed points $\{(\boldsymbol{p}_{i})_{b}\}$ using the Procrustes algorithm. This transformation is then applied to the bone $b$ in the canonical space to derive its corresponding pose in the posed space. This process is repeated for all bones in the skeleton to fit the 3D skeleton to the posed points. We show examples of our fitted skeletons in Fig. 1.
The fitted skeleton can then be used to animate our 3D reconstructions through motion retargeting from existing animations defined on the rigged model used for training. We demonstrate examples of this in Fig. 7.
4.5 Ablation study
We conduct an ablation study to evaluate the impact of different components of our model on its performance and report the results in Tab. 2. Specifically, we examine the advantages of predicting pixel-aligned point maps over directly predicting bone rotations, which are typically used to articulate a rigged mesh model in prior approaches [71, 17, 32]. To test this, we train a version of the method proposed in [71, 32] on our synthetic dataset, where the bone rotation regressor is supervised directly with ground-truth rotations.
The results show that this model performs worse than our proposed approach. We hypothesize that this disparity arises because our representation is more amenable to neural networks. Predicting pixel-aligned point maps effectively transforms the task into a pixel-labeling problem, which is simpler for neural networks to learn. In contrast, directly predicting a chain of 3D rotations for a rigged model is inherently more complex.
Additionally, we investigate the effect of conditioning the predictions of the posed point maps $P$ on the input image features $F$ and the canonical point maps $Q$. When the model is conditioned on both the image features $F$ and the canonical point maps $Q$, it tends to overfit to the features of the training synthetic dataset. In contrast, conditioning the posed point maps $P$ solely on the canonical point maps $Q$ leads to better generalization on out-of-distribution images, as we also demonstrate qualitatively in Fig. 8.
Finally, we analyze the impact of varying the number of layers used in the layered amodal point map predictions. The performance on the Animodel benchmark shows minimal change, as self-occlusions beyond four occluded surfaces are less frequent in the dataset. The number of predicted layers can be easily adjusted to match the complexity of the data.
Figure 8: Effect of different conditioning schemes. Conditioning solely on the canonical point maps $Q$, as opposed to using both the image features $F$ and the canonical point maps $Q$, results in better generalization. This improvement is particularly evident on extreme out-of-distribution images (b).
Method Chamfer Distance (cm) $\downarrow$
Real-Sized Normalized
Predicting bone rotations
(supervised 3D-Fauna [32]) 4.96 ± $\pm\ 1.66$ 1.78 ± $\pm\ 0.59$
Ours - 8-layers 3.13 ± $\pm\ 1.98$ 1.11 ± $\pm\ 0.69$
Ours - 4-layers 3.13 ± $\pm\ 2.13$ 1.11 ± $\pm\ 0.73$
   $P$ conditioned on $F$ and $Q$ 3.29 ± $\pm\ 2.26$ 1.17 ± $\pm\ 0.80$
Table 2: Ablation study on Animodel Horses. We report the bi-directional Chamfer Distance in centimeters (lower is better $\downarrow$) for Animodel Horses (Female subset). The evaluation follows the same protocol as in Tab. 1.
5 Limitations
Figure 9: Typical failure cases. We illustrate representative failure cases caused by (a) extreme viewpoints, (b) shapes and poses far from the training distribution, and (c-d) inaccuracies in the object segmentation masks.
Despite demonstrating surprising generalization, a current limitation of our method is that any additional synthetic 3D models added to the training dataset would have to be in the same canonical space as the training data. Addressing the challenging problem of aligning the canonical spaces of multiple 3D models would allows us to train on significantly larger datasets which could in turn lead to significant gains in performance for our proposed model. Another limitation of our method is that it is not specifically trained to handle occlusions caused by other objects. This is a limitation shared with other methods, such as 3D-Fauna [71, 32]. To address this, we plan to extend the data generation pipeline to include synthetic occlusions. Additionally, as the 3D reconstruction problem is often ambiguous for the unseen parts of objects, our method predicts only the expectation over all possible reconstructions, which can lead to unrealistic results for the invisible regions. We illustrate our typical failure cases in Fig. 9.
6 Conclusion
We presented DualPMs, a novel representation for monocular 3D shape and pose reconstruction of deformable objects. Predicting DualPMs is straightforward for a neural network, and they enable solving a variety of geometric tasks for deformable objects, including not only 3D shape and pose reconstruction but also 3D keypoint localization, skeleton fitting and motion transfer. Despite being trained solely on synthetic data generated from a single horse, our method generalizes effectively to real images, including other categories of quadrupeds, and outperforms state-of-the-art methods on the task of 3D shape and pose reconstruction of horses, cows, and sheep.
Acknowledgments.  We thank Paul Engstler helpful discussions. Ben Kaye is supported by the ESPSRC Centre for Doctoral Training in Autonomous Intelligent Machines and Systems, EP/S024050/1. Tomas Jakab and Andrea Vedaldi are supported by ERC-CoG UNION 101001212. Tomas Jakab is also supported by VisualAI EP/T028572/1.
Appendix
Appendix A Technical details
Network architecture.
We obtain the segmentation mask $M$ using the Segment Anything method [22]. The feature extractor $\Psi$ is based on [84] which combines pre-trained DINOv2 [44] and StableDiffusion [48] networks. Training image features are reduced to a 64-dimensional space using PCA following [70]. The dual point map predictors $\Phi_{Q}$ and $\Phi_{P}$ leverage a convolutional U-Net architecture based on [54], comprising two blocks each and trained from scratch. We predict $N=4$ layers for layered amodal point maps as more have little effect on the performance (Sec. 4.5), likely due to the low frequency of multiple self-occlusions in our datasets. The number of layers can be easily increased should the data require it. The output resolution of the layered point maps is set to 160$\times$160.
Training.
We use the Adam optimizer [21] for training. Our model is trained for 100k steps with a batch size of 12. The learning rate is set to $6\times 10^{-4}$, with a step scheduler applied, featuring a 30k-step period and a decay factor of 0.5.
Training dataset.
The training dataset consists of 33,866 rendered images. We render these images using a single, rigged male horse model featuring two different textures and 50 animated actions, such as running, walking, and drinking. We also randomly sample from a pool of 742 HDRI environmental maps to provide diverse lighting conditions for the training images. We then randomly sample camera viewpoints and poses from the animated actions to generate the training images. Fig. 4 showcases the horse model and some of the generated images used for training.